Here is one question a passenger jet took 3 hours to fly 1,800 miles in the direction of the wind. the return trip against the wind took 4 hours. what the jet's speed in still air and the speed of the wind?

speed of the jet in still air? _____
speed of the wind? _____


Answer by Alan3354(69443)

Speed of jet = j
Speed of wind = w
1800/(j+w) = 3 (took 3 hours with a tailwind)
1800/(j-w) = 4 (took 4 hours with a headwind)
1800 = 3(j+w)
1800 = 4(j-w)
3j + 3w = 1800
4j - 4w = 1800
j + w = 600
j - w = 450
Add the 2 eqns
2j = 1050
j = 525
w = 75